OpenStudy (anonymous):

is y=-9 a solution of y+1=8? show work please

OpenStudy (astrophysics):

So this one should be fairly simple, what you want to do is, substitute -9 for y in your equation and see if it equals the corresponding value of 8. So in mathematical terms, I mean. \[y =-9, ~~~ y + 1 = 8 \implies (-9)+1 = 8\]

OpenStudy (astrophysics):

Now would that be a true statement?

OpenStudy (astrophysics):

You can get rid of the brackets, so -9+1 = 8

OpenStudy (anonymous):

yes? so what about this one then -> m= -1/2 a solution of The inequality m+2<5/4?